You just sold short 450 shares of Wetscope, Inc., a fledgling software firm, at $99 per share. You cover your short when the price hits $89.50 per share one year later. If the company paid $0.81 per share in dividends over this period, what is your rate of return on the investment? Assume an initial margin of 60 percent
Answer #1

Proceeds from short sale / share = Short sale price per share = $ 99

Initial margin = 60%

Initial margin required per share = 60% x $ 99 = $ 59.40

Share purchase price = $ 89.50

Dividend lost per share = $ 0.81

Hence, Dollar Return on short sale = Stock Sale Price – Dividends Paid – Stock Purchase Price = 99 - 0.81 - 89.50 = $ 8.69

Short sale rate of return on investment = Dollar return on short sale / Initial margin = 8.69 / 59.40 = 14.63%